OpenStudy (anonymous):

Hi, I need to establish an identity starting with the left side ln |sec theta| = -ln | cos theta| help?

OpenStudy (anas.p):

Here are the steps, \[\ln (\sec (\theta))=-\ln (\sec (\theta))\]\[e ^{\ln \sec \theta}=e^{-\ln (\sec (\theta))}\]\[\sec(\theta)=1/\cos (\theta)\]\[1/\cos (\theta)=1/\cos (\theta)\]\[1=1\]